OpenStudy (anonymous):

i need someone to please help me with this dba. my mom is going to take all of my privileges away from me if I don't pass...i really don't understand math that well so someone please help.

OpenStudy (anonymous):

Do you know how to find the percent of change? How do you setup the proportion? question 1 of 5

OpenStudy (ahsome):

Don't worry. Legendary @ganeshie8 to the rescue ;)

ganeshie8 (ganeshie8):

\[\large \dfrac{\text{percent change}}{100} = \dfrac{\text{change}}{\text{starting value}} \]

ganeshie8 (ganeshie8):

thats the proportion^^ multuply both sides by 100 to get the `percent change`

ganeshie8 (ganeshie8):

Notice that we're taking the percent in reference to the "starting value"

OpenStudy (anonymous):

ive got another question

ganeshie8 (ganeshie8):

Okay, ask... watch this for more info on calculating percent of change and examples http://www.virtualnerd.com/pre-algebra/percents/change/change-examples/percent-change-definition

OpenStudy (anonymous):

what is the percentage proportion formula? What is the caution about for this formula? Question 2 of 5

ganeshie8 (ganeshie8):

above given is the percentage proportion formula

OpenStudy (anonymous):

well what is the caution then?

ganeshie8 (ganeshie8):

Make sure you divide the "original/starting value" and not the "new/final" value Thats the caution they're asking about I think

OpenStudy (anonymous):

here's another question

OpenStudy (anonymous):

@ganeshie8

ganeshie8 (ganeshie8):

yeah ?

OpenStudy (anonymous):

what is a ratio and how can they be compared?

ganeshie8 (ganeshie8):

http://www.mathsisfun.com/numbers/ratio.html

OpenStudy (anonymous):

@ganeshie8 help plzzz

ganeshie8 (ganeshie8):

two ratios can be compared by forming a proportion

OpenStudy (anonymous):

ganeshie plzz sirrr

ganeshie8 (ganeshie8):

pm me surd :)

OpenStudy (anonymous):

what are the three different ways to express a ratio?

OpenStudy (anonymous):

@ganeshie8

OpenStudy (anonymous):

Like this, 3:5 3/5 and 3 to 5 (the "to" is important

OpenStudy (anonymous):

thanks one more?

OpenStudy (anonymous):

Sure

OpenStudy (anonymous):

hold on one sec

OpenStudy (anonymous):

what is the simple interest equation and how do you solve it

OpenStudy (anonymous):

Let's use this. Equation: A = P(1 + rt) Calculation: First, converting R percent to r a decimal r = R/100 = 3.875%/100 = 0.03875 per year, then, solving our equation A = 10000(1 + (0.03875 × 5)) = 11937.5 A = $ 11,937.50 The total amount accrued, principal plus interest, from simple interest on a principal of $ 10,000.00 at a rate of 3.875% per year for 5 years is $ 11,937.50.

OpenStudy (anonymous):

i don't get it

OpenStudy (anonymous):

walk me through it

OpenStudy (anonymous):

Okay, hold on a second.

OpenStudy (anonymous):

So, let's say we have $4500.00 at 9.5% for 6 years. You can order it like this So, first you take the percent and divide it by one hundred. 9.5 divided by 100 = 0.095. Now you order it like this (4500) (0.095) (6) Now just solve. So, 4500 times 0.095 times 6 = $2565.

OpenStudy (anonymous):

We can do another if you want.

OpenStudy (anonymous):

hold on don't leave i want to figure this out

OpenStudy (anonymous):

I won't.

OpenStudy (anonymous):

yeah these are questions that are going to be asked during a dba and the review of it says "know the simple interest equation and be able to use it in solving a problem

OpenStudy (anonymous):

Yeah, that is it. Just divide the percent by 100, then multiply everything.

OpenStudy (anonymous):

i am not going to do good on this dba

OpenStudy (anonymous):

Just do you best.

OpenStudy (anonymous):

i cant remember anyhting and i can't talk to my teachers

OpenStudy (anonymous):

but if i dont get this question i am sure im going to get the other four right and that is four out of five so...

OpenStudy (anonymous):

can you walk me through how to divide please just to make sure i got it down

OpenStudy (anonymous):

Yes.

OpenStudy (anonymous):

So, I think long division is easiest.

OpenStudy (anonymous):

Use this. http://www.mathsisfun.com/long_division.html

OpenStudy (anonymous):

i might just do it on a calculator when im on the phone with her

OpenStudy (anonymous):

Yes, that might work too.